OpenStudy (anonymous):

Find the zeros of the polynomial function. f(x) = x^3 - 6x^2 + 9x

OpenStudy (anonymous):

\[ x^3 - 6x^2 + 9x =x( x^2 - 6 x +9) \] Can you finish it?

OpenStudy (anonymous):

\[ x^3 - 6x^2 + 9x =x( x^2 - 6 x +9)= x(x-3)^2 \]

OpenStudy (anonymous):

What are the zeros now?

OpenStudy (anonymous):

I can finish it

OpenStudy (anonymous):

the zeros are x=0 x=3 x=3

OpenStudy (anonymous):

A) x = 0, x = 3

OpenStudy (anonymous):

The answer was A) x = 0, x = 3. Thanks for your help @eliassaab

OpenStudy (anonymous):

3 is double root/